OpenStudy (anonymous):

Find the exact values of all solutions in the interval 0=/< x<2pi for sin [x+(pi/4)] + sin [x-(pi/4)] = -1

OpenStudy (anonymous):

do you know sin(a+b) and sin(a-b) formulae?

OpenStudy (anonymous):

the product to sum formula?

OpenStudy (anonymous):

or sum and difference?

OpenStudy (anonymous):

sin(a+b)=sin a cos b+sin b cos a

OpenStudy (anonymous):

oh yeah! i know that

OpenStudy (anonymous):

sin(a-b)=sin a cosb-sin b cos a so if you add sin(a+b) and sin(a-b) you're left with 2 sina cos b here a=x,b=pi/4 so 2sina/sqrt(2)=-1 sin a=-1/sqrt(2)

OpenStudy (anonymous):

so you just use that formula to substitute? oh thanks! you're very helpful

OpenStudy (anonymous):

yeah no problem.

OpenStudy (anonymous):

sinxcospi/4 - cosxsinpi/4

OpenStudy (anonymous):

sinxcospi/4 - cosxsinpi/4 = -1 what do i do next?

OpenStudy (anonymous):

Expand both you have expanded only one two terms will cancel. You'll get sin x cos pi/4=-1 cos pi/4=sqrt(2)